OpenStudy (anonymous):

A sculpture was purchased for $4,000. The value of the sculpture increases at a rate of 4% year. How would I figout how much the sculpture would be worth after 10 years?

OpenStudy (anonymous):

4000(1+4/100)^10 Now solve

OpenStudy (anonymous):

About $5,921?

OpenStudy (amistre64):

A{0} + A{0}(.04) A{0} (1.04) A{n} = A{n-1}(1.04) = (A{n-2}(1.04))(1.04) = ((A{n-3}(1.04))(1.04))(1.04) = (A{n-r} (1.04)^r when A{n-r} = A{0} when r=n A{n} = (A{n-n} (1.04)^n = A{0} (1.04)^n A{0} = 4000 so... A{n} = 4000(1.04^n); when n=10 is the value you want right?

OpenStudy (anonymous):

Yeap

OpenStudy (anonymous):

Whoa. Um, yes. So my calculation was correct?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Is amistre64 just breaking it down for me?

OpenStudy (amistre64):

Yes, the formula comes from the recurssion

OpenStudy (anonymous):

Thank you so much guys! :)

OpenStudy (anonymous):

amistre Let's give him a break with the derivation, shall we? :-)

OpenStudy (amistre64):

but..but... i need the practice lol

OpenStudy (anonymous):

In how many years will the sculpture be worth $10,000? Round to the nearest tenth of a year. :/

OpenStudy (anonymous):

10000=4000(1+4/100)^T

OpenStudy (anonymous):

Now solve in the same way I helped u in the previous question

OpenStudy (anonymous):

4160?

OpenStudy (anonymous):

I don't think so, check it again

OpenStudy (anonymous):

4000?

OpenStudy (anonymous):

The sculpture was initially purchased for 4000 ,so

OpenStudy (anonymous):

:/ I'm stuck. After 50 years?

OpenStudy (amistre64):

lets view it like this: A = P (1+r)^t ; and we solve for t, divide out the P A/P = (1+r)^t ; now we go to logs to get the "t" ln(A/P) = [ln(1+r)^t] ln(A/P) = t ln(1+r) ; divide out the ln(1+r) ln(A/P) ------ = t ln(1+r)

OpenStudy (anonymous):

Wait, what does A, P, and r stand for again?

OpenStudy (anonymous):

Hello?